### v2

# ExMAG: Learning of Maximally Ancestral Graphs

Abstract
As one transitions from statistical to causal learning, one is seeking the most appropriate causal model. Dynamic Bayesian networks are a popular model, where a weighted directed acyclic graph represents the causal relationships. Stochastic processes are represented by its vertices, and weighted oriented edges suggest the strength of the causal relationships. When there are confounders, one would like to utilize both oriented edges (when the direction of causality is clear) and edges that are not oriented (when there is a confounder or not a relationship), yielding mixed graphs. A little-studied extension of acyclicity to this mixed-graph setting is known as maximally ancestral graphs with consideration of confounders.
We propose a score-based learning algorithm for learning maximally ancestral graphs. A mixed-integer quadratic program is formulated, and an algorithmic approach is proposed, in which the pre-generation of exponentially many constraints is avoided by generating only violated constraints in the so-called branch-and-cut (“lazy constraint”) method. Comparing the novel approach to the state-of-the-art, we show that the proposed approach turns out to produce more accurate results when applied to small and medium-sized synthetic instances containing up to 25 variables.
1 Introduction
As one transitions from statistical to causal learning [47], one is seeking the most appropriate causal model. Dynamic Bayesian networks (DBN) [17, 35] are a popular model, where a weighted directed acyclic graph represents the causal relationships. Stochastic processes are represented by its vertices, and weighted oriented edges suggest the strength of the causal relationships. The key challenge in learning DBNs is confounding.
To illustrate the challenge of confounding, let us consider Simpson’s paradox. Simpson’s paradox shows that without considering confounding factors in statistical analysis [34], the direction of causality can be mis-estimated completely. An textbook example [34] comes from the Berkeley graduate admissions [3]. The data show that women find it harder to get admitted to Berkeley graduate schools. Nevertheless, this is because women tend to apply to departments that have lower admission rates. In this example, the choice of the graduate school is the confounder, impacting the probability of admission. Confounding is prevalent throughout high-dimensional statistics [32, 21, e.g.], such as in biomedical sciences.
Specifically, in biomedical sciences, confounders such as socio-economic status, age, or lifestyle factors can distort the true causal relationship between treatments and outcomes [52]. Techniques such as instrumental variables [39, 29], propensity score matching [43], and double machine learning [11] have been widely used to mitigate the effects of confounding in clinical trials and observational studies. In bioinformatics, particularly in genome-wide association studies (GWAS), confounders, including population stratification and environmental exposures, must be controlled to avoid biased estimates in genetic association studies [9, 22]. To mitigate such biases, statistical models that explicitly account for hidden confounders, such as spectral methods and latent variable models, are often employed [22]. Furthermore, meta-analysis and sensitivity analysis are often employed to evaluate the robustness of findings in the presence of potential confounders, especially when combining results from multiple studies [6, 33]. These methodologies ensure that the conclusions drawn are reliable and actionable, improving the credibility of statistical models across disciplines.
In statistical theory, [8] study confounding in detail, and many subsequent works develop further methods. [31] show that leveraging the dominant eigenstructure of time series may improve performance of estimation. Anchor regression, for instance, bridges the gap between causality and robustness by addressing heterogeneity in data [44]. Other significant contributions include spectral deconfounding, a technique designed to mitigate the effects of hidden confounders in high-dimensional settings [7]. This approach provides a framework for robust predictions in the presence of shifts in data distributions. Similarly, the invariance principle has emerged as a cornerstone of causal inference, linking causal structure to robust statistical models [6]. Furthermore, the concept of doubly robust inference offers an alternative framework for addressing hidden confounding factors, combining model robustness with efficiency in high-dimensional scenarios [22]. Together, these developments represent a significant step forward in understanding and addressing the challenges posed by complex causal systems.
While there is a long history of the study of confounding, as suggested above, the extensions of DBNs to allow for confounding are rather more recent. Instead of estimating a Directed Acyclic Graph (DAG), one could estimate a Maximal Ancestral Graph (MAG), cf. [41]. These allow for both direct and indirect relationships among variables modelled as directed and undirected edges, even in the presence of confounding factors. In particular, MAGs can represent feedback loops and bidirectional relationships that DAG-based models, such as DBNs, cannot. This makes MAGs a more powerful tool for capturing the complex dynamics of real-world causal systems.
There are only a few studies of MAG estimation [10, 38, 12, 24, 27, 26, 15]. [40, 49] are applicable to the both discrete and nonparametric cases, which extend DAG to MAG, ADMG or Bow-Free Acyclic Path Diagrams, see more definition in 2. Factorization in MAG is not directly decomposable into individual variables and their parent sets, as in DAGs, but must instead consider components connected by bidirected paths (termed districts or c-components), cf. [40], although [12] proposed to use Markov equivalence classes (MEC) instead. In 2021, [10] introduced a first mixed-integer programming (MIP) formulation, but the number of variables scales with the number of c-components, i. e., exponentially with the number of vertices in the worst case. Such formulations are also known as extended formulations [13]. The same year, [38] explored a constraint-based approach for MAG discovery, leveraging conditional independence testing. Additionally, [53] addresses exogenous covariates in causal formulation that helps explain the heterogeneity in both sampling and causal mechanisms. The dissertation of [24] presented an extension of the imsets of [48] from directed acyclic graphs (DAGs) to towards MAGs [27], which allows for the use of the methods of Studený, and a score-based heuristic [26]. More recently, the paper by [15] enhanced the scalability of methods of [10] by utilizing linear programming (LP) relaxations instead of solving the MIP.
Our approach proposes the first compact formulation of MAG estimation within Mixed-Integer Nonlinear Programming (MINLP), in contrast with the so-called extended formulations of [10, 15]. While both the extended and compact formulations ensure that confounding factors are properly accounted for and the true underlying data-generating process is better represented by the model, our implementation scales further, from 4-5 stochastic processes in the extended formulation to 25 or more stochastic processes with the proposed compact formulation.
1.1 Motivating Example
Let us revisit the Berkeley graduate admission paradox example of page 1. As in most paradoxes, there is no violation of logic in Simpson’s paradox, just a violation of of intuition. The poor intuition being violated in this case is that a positive association in the entire population should also hold within each department. Overall, females in these data did have a harder time getting admitted to graduate school. But that arose, because female applicants chose the departments that were the most difficult to gain admission to for anyone, male or female. In this example, gender influences the choice of department, and the department influences the chance of admission. Controlling for department reveals a more plausible direct causal influence of gender. If one represents gender as $G$, department as $D$, acceptance as $A$, and academic ability as $U$, the potential confounder, the DAG form is
[…]
Our method, ExMAG, is able to reveal the confounders in this Berkeley graduate admission example, as illustrated in the notebook in supplementary materials and Figure 1.
Figure 1: Performance of ExDBN of [45] (top) and ExMAG (bottom) on the Berkeley graduate admission example. While the DBN identified by ExDBN suggests a causal relationship between gender and admission, ExMAG correctly identifies the confounding.
2 Graphs and Properties
Maximal ancestral graphs (MAGs), first introduced by [41], provide a framework for modeling distributions through conditional independence (CI) relations. Compared with directed acyclic graphs (DAGs), MAGs allow for latent confounders, accommodating data that arises from distributions with more complex independence structures and revealing hidden states in the graphs. While DAGs allow for the efficient computation of maximum likelihood estimates (MLEs) and scoring (e.g., via BIC), these properties are challenging to extend to MAG due to their structural and computational complexity [25].
Directed Acyclic Graph
A DAG is a directed graph $G=(V,E)$ such that there are no directed cycles. That is, there is no sequence of distinct vertices $v_{1},v_{2},\dots,v_{k}\in V$ such that $(v_{i},v_{i+1})\in E$ for all $11\leq i\leq k-$ 1 and $(v_{k},v_{1})\in E$.
Bow-Free Acyclic Path Diagrams
Let $G$ be a bow-free ADMG (Acyclic Directed Mixed Graph). The graph $G$ is bow-free if and only if there does not exist a directed and bidirected edge between a pair of adjacent vertices. More formally, for any $v_{i},v_{j}\in V$, there cannot exist both $v_{i}\to v_{j}$ (or $v_{i}\leftarrow v_{j}$) and $v_{i}\leftrightarrow v_{j}$ simultaneously, as described by [50] .
ADMG
Mixed graphs feature two types of edges: directed ($\rightarrow$) and bidirected ($\leftrightarrow$). Mixed graph $\mathcal{G}$ thus consists of a vertex set $\mathcal{V}$, a set of directed edges $\mathcal{E}$ and undirected edges $\mathcal{U}$, where $\mathcal{E}$ are (ordered) pairs of distinct vertices, while $\mathcal{U}$ are (unordered) 2-element subsets of vertices. For a directed edge in $\mathcal{E}$ connecting vertices $v$ and $w$, we say these two vertices are the endpoints of the edge and the two vertices are adjacent (if there is no edge between $v$ and $w$, they are non-adjacent). For a vertex $v$ in $\mathcal{V}$, we define the parents, spouses, ancestors, and district of $v$, respectively as:
$\text{pa}_{\mathcal{G}}(v)=\{w:w\to v\text{ in }\mathcal{G}\},$
$\text{sp}_{\mathcal{G}}(v)=\{w:w\leftrightarrow v\text{ in }\mathcal{G}\}$
$\text{an}_{\mathcal{G}}(v)=\{w:w\to\cdots\to v\text{ in }\mathcal{G}\text{ or }w=v\},$
$\text{dis}_{\mathcal{G}}(v)=\{w:w\leftrightarrow\cdots\leftrightarrow v\text{ in }\mathcal{G}\text{ or }w=v\}.$
Given a directed mixed graph $\mathcal{G}$, the districts define a set of equivalence classes of nodes in $\mathcal{G}$. The district for node $v$ is defined as the connected component of $v$ in the subgraph of $\mathcal{G}$ induced by all bidirected edges. As in a DAG, a mixed graph $\mathcal{G}$ is acyclic if it contains no directed cycles in $\mathcal{E}_{d}$, i.e., an acyclic directed mixed graph (ADMG) [24].
A directed mixed graph G is called an ancestral ADMG if the following condition holds for all pairs of nodes $v$ and $w$ in $\mathcal{G}$:
$\text{If }v\neq w\text{ and }w\in\text{an}_{\mathcal{G}}(v)\cup\text{sp}_{\mathcal{G}}(v),\text{ then }v\notin\text{an}_{\mathcal{G}}(w),$
which is written as, $\mathcal{G}$ is an ancestral ADMG if it contains no directed cycles ($v\to u\to\ldots\to w\to v$) or almost directed cycles [10, 24].
An almost directed cycle is of the form $v\to u\to\ldots\to w\leftrightarrow v$; in other words, $\{v,w\}\in\mathcal{E}_{w}$ is a bidirected edge, and $v\in\text{an}_{\mathcal{G}}(w)$ [10].
Inducing Paths
An inducing path from a variable $X$ to a variable $Y$ in a directed graph $G=(V,E)$ is a path $P=(X=v_{0},v_{1},\dots,v_{n}=Y)$ such that for all intermediate nodes $v_{i}$ (where $11\leq i\leq n-$ 1), $v_{i}\in Z$, where $Z$ is the conditioning set. If the path is blocked by conditioning on $Z$, then it is an inducing path.
A node j on a non-overlapping path is called a collider if contains a non-overlapping subpath (i; j; k) with two arrowheads into j. In mathematical form, collider is represented as
$\text{collider}_{G}(u)=\left\{u:\exists w,v:\begin{array}[]{c}w\to u\leftarrow v\vee\\ w\leftrightarrow u\leftarrow v\vee\\ w\to u\leftrightarrow v\vee\\ w\leftrightarrow u\leftrightarrow v\end{array}\right\}.$
m-separation
For an ADMG $\mathcal{G}$, given a subset $W\subseteq\mathcal{V}$, the induced subgraph $\mathcal{G}_{W}$ is defined as the graph with vertex set $W$ and edges in $\mathcal{G}$ whose endpoints are both in $W$. Also for the district of a vertex $v$ in an induced subgraph $\mathcal{G}_{W}$, we may denote it by $\text{dis}_{W}(v)=\text{dis}_{G_{W}}(v)$.
Graphs are associated with conditional independence relations via a separation criterion; in the case of ADMGs, we use m-separation. Graphs encode conditional independence relations through separation criteria. For ADMGs, m-separation is used, which generalizes d-separation for DAGs to account for bidirected edges. A path between two vertices $a$ and $b$ is m-connecting given a conditioning set $C\subseteq\mathcal{V}$ if:
  1. 1.
    $a$ and $b$ are endpoints of the path;
  2. 2.
    Every noncollider on the path is not in $C$;
  3. 3.
    Every collider is in $\text{an}_{G}(C)$.
Graphs that encode the same set of CI relations are said to belong to the same MEC. This equivalence class structure is critical for reducing redundancy during graph exploration and scoring.
Maximal Ancestral Graph
An ADMG $\mathcal{G}$ is called a maximal ancestral graph (MAG) if:
  1. (i)
    For every pair of nonadjacent vertices $a$ and $b$, there exists some set $C$ such that $a,b$ are m-separated given $C$ in $\mathcal{G}$ (Maximality);
  2. (ii)
    For every $v\in\mathcal{V}$, $\text{sib}_{\mathcal{G}}(v)\cap\text{anc}_{\mathcal{G}}(v)=\emptyset$ (Ancestrality).
We refer to [24] for multiple examples.
3 Brief Introduction To Mixed Integer Quadratic Programming
Let us also provide a short introduction to mixed-integer quadratic programming. An optimization problem is called a mixed-integer quadratically constrained quadratic program (MIQCQP) if it is of the form
$\underset{x\in\mathbb{R}^{n}}{\min}\quad$ $x^{T}Qx+q^{T}x,$ (1)
s.t. $x^{T}Q_{i}x+q_{i}^{T}x\leq a_{i},$ (2)
$Ax\leq b,$ (3)
$x\in F$ (4)
where $Q,Q_{i}\in\mathbb{R}^{n,n}$, $q,q_{i}\in\mathbb{R}^{n}$, $A\in\mathbb{R}^{m,n}$, $a\in\mathbb{R}^{k}$, $b\in\mathbb{R}^{m}$, $F$ is a product of the form
$F=\underbrace{\mathbb{R}\times\ldots\times\mathbb{R}}_{n-r\text{ times}}\times\underbrace{\mathbb{N}\times\ldots\times\mathbb{N}}_{r\text{ times}}$ (5)
and $m,n,k,r\in\mathbb{N}$. In the mathematical form, (1) is often called the cost or loss function, (2) represents the quadratic constraints, (3) is the linear constraints, and $F$ is the set that enforces the integrality constraints for the $r$ components of the decision variable $x$.
Mixed-integer quadratic programs have been shown to be NP-hard [18], which often leads to an exhaustive demand for computational resources. The algorithms used to solve MIQP are typically branch-and-bound or cutting plane [14, 5, 51, 30]. Both of these algorithmic treatments are often employed together, often with the addition of a presolving step, the use of heuristics, and parallelism. The aforementioned allows many modern solvers to solve even large problems despite the NP-hardness. Some of these solvers are open source (like SCIP and GLPK), and others are commercial (GUROBI and CPLEX). The powerful infrastructure present in these solvers can be made use of together with additional problem-specific modifications to deliver high-quality solutions.
Due to the exhaustive nature of the algorithms mentioned in the previous paragraph, global convergence is guaranteed [2]. Furthermore, convergence to the global solution may be tracked, and the error estimated by computing the dual problem of (14). The dual of the problem is then used to compute the so-called MIP GAP as follows
MIP $\text{MIP GAP}=\frac{\left|J\left(x^{*}\right)-J_{\text{dual}}\left(y^{*}\right)\right|}{\left|J\left(x^{*}\right)\right|},$ (6)
where $x^{*}$ and $y^{*}$ are the current best solutions of the primal and dual problems, respectively, and $J$ and $J^{*}$ are the cost functions of the primal and dual problems, respectively. The MIP GAP ensures that we can assess the quality of the minimization during solution time and terminate the computation when the result is good enough (small enough MIP GAP). Furthermore, if the gap reaches 0 at any point, we are sure that the current solution is a global optimum.
4 Formulation of the Mixed Integer Quadratic Program
Recent works on high-dimensional confounding or deconfounding clarify the connections between distributional robustness, replicability, and causal inference [44, 22]. Distributional robustness differs significantly from traditional robust statistical methods [28, 23], which typically handle outliers in the training data, while our work focuses on evaluating the existence of a confounding factor.
In this section, we inherit from distributional robustness and present the formulation of the Mixed Integer Quadratic Program (MIQP) to infer the causal structure. Since we cannot observe all relevant variables, we must deal with the situation of hidden confounding. The problem is formalized in the following form corresponding to a structural equation model (SEM) [4, 36]:
$Y_{i}\leftarrow X_{i}\tilde{w}+g(H_{i},A_{i})+\epsilon_{Y,i},$
where:
  • $\epsilon_{Y,i}$ is the noise term, independent of all variables that appear ”upstream” from $Y_{i}$.
  • $A_{i}$ is an exogenous variable, though not considered in the following sections.
  • $\leftarrow$ represents an algebraic equality, implying a structural causal relationship.
  • $H_{i}$ is the unobserved confounding variable vector.
If non-zero components of the vector $\tilde{w}$ are correlated with certain components of $X_{i}$, these components are defined as causal $X$-variables for $Y_{i}$. This means:
[…]
For the scenario that there are no exogenous variable perturbations, we describe this with an additive noise model with hidden states as follows:
$Y_{i}$ $\leftarrow X_{i}\tilde{w}+H_{i}\theta+\epsilon_{Y,i},$ (8)
$X_{i}$ $\leftarrow H_{i}\gamma+\epsilon_{X,i},$ (9)
where $\epsilon_{X,i}$, $\epsilon_{Y,i}$ and $H_{i}$ are mutually independent. $X_{i}$ is the observed $p\times 1$ covariate vector. $H_{i}$ is the unobserved $q\times 1$ confounding variable vector. Our goal is to infer the confounding-free regression parameter $\tilde{w}$ and stabilize the prediction of the relationship between $Y$ and $X$.
4.1 Connecting to Causality
The causal parameter $\tilde{w}$ can be seen as minimizing the worst-case risk:
$\tilde{w}=\arg\min_{w}\max_{P\in\mathcal{P}}\mathbb{E}_{P}\left[(Y_{i}-X_{i}w)^{2}\right],$
where $\mathcal{P}$ is a class of distributions containing perturbations of the original distribution, including confoundings. This modeling highlights the inherent connection between causality and distributional robustness [16, 37, 42].
In this perspective, we present the formulation of the Mixed Integer Quadratic Program (MIQP) used to infer the causal structure, with a new binary matrix $B=[b_{j,k}]\in\{0,1\}^{d\times d}$ introduced to account for relationships explained by confounding factors, alongside the weight matrix $W$ and binary adjacency matrix $E$ adopted from the ExDAG model [46]. Weight matrix $W$ represents the model weights. Whenever an entry in $W_{j,k}$ is non-zero, the respective value in either $E_{j,k}$ (directed edge) or $B_{j,k}$ is nonzero (bidirected edge). At the same time, we extend the existing formulation by introducing an additional binary input parameter $f_{j,k}$ for each pair of variables $(j,k)$, where $j\neq k$, indicating that there is no direct causal relationship between variables $j$ and $k$, but $j$ and $k$ might have a common cofactor. This follows from the meaning of the edges in a MAG $\rightarrow$ edge implies a direct causal relationship but does not rule out a possible latent confounding, $\leftrightarrow$ means no direct causal relationship. Formally, we define the Directed Edge Matrix $E$ as
$e_{j,k}=\begin{cases}1,&\text{if $j\to k$},\\ 0,&\text{otherwise},\end{cases}$ (10)
and the Bidirected Edge Matrix $B$ as
$b_{j,k}=\begin{cases}1,&\text{if $j\leftrightarrow k$},\\ 0,&\text{otherwise}.\end{cases}$ (11)
The new input matrix $F$, is by the definition,
$f_{j,k}=\begin{cases}1,&\mbox{if }j\not\to k\quad\mbox{and}\quad k\not\to j,\\ 0,&\mbox{otherwise}.\end{cases}$ (12)
Based on these assumptions, one can substitute algebraic equalities 9 into 8, and the work of [46] is extended as explained in the next section.
4.2 MILP Formulation
The cost function for the Mixed Integer Quadratic Program of ExMAG is the following $l_{q}$ norm:
$l_{q}$ $=\sum_{i=1}^{n}\sum_{j=1}^{d}\left[X_{i,j}-\sum_{k=0;k\neq j}^{d}X_{i,k}w_{k,j}\right]^{q}$ (13)
$\quad+\lambda\sum_{j=0}^{d}\sum_{k=0}^{d}(e_{j,k}+b_{j,k}),$
where:
  • $X_{i,j}$ represents the value of the $j$-th variable for the $i$-th data point;
  • $w_{k,j}$ represents the weight of the edge from variable $k$ to variable $j$;
  • $e_{j,k}$ is the binary decision variable indicating the presence of a directed edge from $j$ to $k$;
  • $b_{j,k}$ is the binary decision variable indicating the presence of a bidirected edge between $j$ and $k$;
  • $\lambda$ is a regularization parameter controlling the trade-off between the model fit and the edge penalty.
  • The exponent $q\in\mathbb{N}$ can take values $q=1$ or $q=2$. While setting $q=1$ leads to a mixed-integer linear program (MILP), which case would not correspond to the maximum likelihood estimator. If we choose $q=2$, resulting in a mixed-integer quadratic problem(MIQP).
As in the ExDAG model, the weights are bounded by introducing a large constant $c$ that ensures the weights are appropriately constrained when there is an edge between the variables. The bounding avoids bilinear terms in the cost function in (13). As a matrix form, the weight constraints can be restrained as:
$-c\cdot(E+B)$ $\leq W\leq c\cdot(E+B)$ (Weight Constraint)
$E+B$ $\leq\mathbf{1}$
The (4.2) means that there cannot be a directed as well as a bidirected edge between the same two vertices. Additionally, we enforce that the bidirected matrix is symmetric, i.e.,
$B=B^{T}.$ (14)
If $f_{j,k}=1$, then we know there is no direct causal relationship between $j$ and $k$, and therefore, $e_{j,k}=0$. Formally, this translates to condition
$F+E\leq\mathbf{1}.$ (15)
Inversely, $f_{j,k}=0$ implies a directed edge rather than a bidirected edge between $j$ and $k$. Therefore,
$B\leq F.$ (16)
Lastly, we must enforce conditions for directed or almost directed cycles and inducing paths. Those conditions are enforced lazily using a separation routine explained later. Directed cycles are enforced in a way adopted from [46]. Therefore, they are left out of this paper. Suppose an almost directed cycle is formed by edges in set $E$ and a bidirected edge $(u,v)$. Then, this almost directed cycle is forbidden by constraint
$b_{u,v}+\sum_{(j,k)\in E}e_{j,k}\leq|E|.$
Similarly, if there is an inducing path formed by path $P$ that contains bidirected edges, and set $E$ contains all directed edges that participate in the ancestor relationship (including multiple paths) between the inner points of the path and the terminals of $P$, this inducing path is forbidden by
$\sum_{(j,k)\in P}b_{j,k}+\sum_{(j,k)\in E}e_{j,k}\leq|E|+|P|-1.$
Note that the second condition does not necessarily eliminate the inducing path, as the optimizer might forbid one of the directed edges in $E$ without influencing the ancestor relationship. This results in the path $P$ being found in the next iteration, with a smaller set of directed edges, and the process is repeated.
By enforcing these constraints, we ensure that the MIQP correctly models the causal relationships between the variables while respecting the independence structure defined by $f_{j,k}$ and the potential confounding relationships captured by $b_{j,k}$.
5 Separation Routine for the Maximal Ancestral Graphs
The main contribution of this section is the separation routine that allows us to identify whenever a graph is not an instance of a maximal ancestral graph. To do so, we need to identify directed cycles, almost directed cycles, and inducing paths. The directed cycles can be found in $\mathcal{O}(d^{2})$ using depth-first-search (DFS); such an approach can be found in [46]. For both inducing paths and almost directed cycles, we will use a distance matrix constructed on the graph of directed edges. Such distance matrix can be obtained using the Floyd-Warshall algorithm [20].
Once having the distance matrix, to check for almost directed cycles, we can iterate over all bidirected edges and test whether the distance between the endpoints is finite, i.e., we have a directed path connected by a bidirected edge. See Algorithm 1 for details.
Input: directed edges $e$, bidirected edges $b$
function Almost-Directed-Cycles($e$, $b$)
     $d\leftarrow$ Distance-Matrix($e$)
     for all $j\in{1,2,\ldots,d}$ do
         for all $k\in{1,2,\ldots,d}$ do
              if $b_{j,k}==1\And d_{j,k}<\infty$ then
                  $E=$ Trace-Distance-Matrix($d$,$e$,$j$,$k$)
                  Found cycle formed by edges $E$ and $j\leftrightarrow k$
              end if
         end for
     end for
end function
Algorithm 1 Function to identify almost directed cycles.
In the case of inducing paths, the problem is more complicated. We use DFS started from each of the vertices. Once started from vertex $x$, this DFS routine checks for all possible inducing paths that terminate in $x$. For efficiency, a set of all possible endpoints of the path is held. Once this set is empty, the DFS search is terminated, and no further exploration is performed. The set is updated using the distance matrix calculated from the graph of the directed edges. If we consider a vertex $y$, $x$ must be either its ancestor (meaning that possible endpoints for $y$ remain unchanged) or it is among the points that are reachable from $y$ (meaning that the possible endpoints for $y$ are replaced with their intersection with the set of all points reachable from $y$). See Algorithm 2 for details.
Input: directed edges $e$, bidirected edges $b$
function Inducing-Paths($e$, $b$)
     $d\leftarrow$ Distance-Matrix($e$)
     for all $j\in{1,2,\ldots,d}$ do
         Inducing-Paths-DFS($d$, $e$, $b$, $j$, $j$, $\{1,2,\ldots d\}$, $[j]$)
     end for
end function
function Inducing-Paths-DFS($d$, $e$, $b$, start, $u$, possible endpoints, path)
     if possible endpoints are empty then return
     end if
     if Len(path) $>2\And u$ in possible endpoints  then
         Found-Inducing-Path($d$,$e$, path)
     end if
     for all $v\in{1,2,\ldots,d}$ such that $e_{u,v}=1$ do
         if $d_{v,start}<\infty$ then v-endpoints = possible endpoints
         elsev-endpoints = possible endpoints $\cap\{x\mid d_{v,x}<\infty\}$
         end if
         Inducing-Paths-DFS($d$,$e$,$b$,start,$v$,v-endpoints,path$+v$)
     end for
end function
function Found-Inducing-Path($d$,$e$, $P$)
[…]
    for all vertices $j\in$ $P$ and $j\not\in\{P_{0},P_{|P|}\}$ do
        $E=E\cup$ Trace-Distance-Matrix($d$,$e$,$j$,$P_{0}$)
        $E=E\cup$ Trace-Distance-Matrix($d$,$e$,$j$,$P_{|P|}$)
    end for
    Found inducing path formed by path $P$ and directed edges $E$
end function
Algorithm 2 Function that identifies inducing paths.
Once having the bidirected edges in the inducing paths and almost directed cycles, we need to trace back the Floyd-Warshall distance matrix to find all edges that form the cycle or the ancestor relationship. This is done using calls to the function in Algorithm 3.
If directed cycles, almost directed cycles, and inducing paths are found, the algorithm applies lazy constraints in 4.2 and in 4.2. Otherwise, we know that the program converged to the optimum, and we have a maximal ancestral graph, which minimizes (13).
Input: distances $d$ defined by directed edges $e$, start point $j$, and endpoint $k$
Output: edges on any path from $j$ to $k$
function Trace-Distance-Matrix($d$, $e$, $j$, $k$)
     if $d_{j,k}==\infty$ then return
     end if
     visited = $\{(j,k)\}$
     stack = stack with $(j,k)$
[…]
     while stack is not empty do
         $u,v\leftarrow$ Pop(stack)
         for all $w\in{1,2,\ldots,d}$ s t. $d_{u,w}+d_{w,v}<\infty$ do
              visited $\leftarrow$ visited $\cup\{(u,w),(w,v)\}$
              if $e_{u,w}$ then edges $\leftarrow$ edges $\cup\{(u,w)\}$
              else if $e_{w,v}$ then edges $\leftarrow$ edges $\cup\{(w,v)\}$
              end if
              add $\{(u,w),(w,v)\}$ to stack
         end for
     end while
end function
Algorithm 3 Functions that help in the separation routine.
6 Experimental Evaluation
Figure 2: SHD values (in the vertical axis) for different settings of $d$ (in the horizontal axis) and $n$ (horizontal choice of the graph). The plots in the vertical dimension differ according to the dataset used. The channel shows the variance of the results for different choices of seed. See supplementary materials for results on more datasets.
6.1 Used Datasets
We test the ExMAG algorithm on both synthetic and real-world datasets. The first synthetic dataset is based on the Erdös-Rényi model (ER) [19], in which the ground truth graph is randomly selected from all graphs with $d$ vertices and $m$ edges (parameter of the experiment, for example, dataset ER-5 contains $5$ edges per variable, i.e., $m=5\cdot d$). The weights of the graph are randomly sampled form the set $(-2.0,-0.5)\cup(0.5,2.0)$.
Once the ground truth model is created, then the training data are generated using the structural model equation 8,  9 ($H_{i}$ set to 00). Then, $20\,\%$ of variables are treated as latent variables and hidden from the training data. The respective columns and rows from the ground truth weight matrix $W$ have also been removed. Finally, $20\,\%$ of edges between variables not connected by an edge in the ground truth data are marked in $F$.
The second dataset uses randomly generated bow-free (BF) graphs, a subset of all possible MAGs. A bow-free graph is defined as a graph that does not contain a pair of vertices $i,j$, such that $i\rightarrow j$ and at the same time $i\leftrightarrow j$. The BF graph generation process has two parameters. First, the probability of a directed edge. Second, the probability of bidirected edge. The generation process is as follows. First, a bow-free graph with the given edge probabilities is generated randomly. The weights of the sampled graphs are randomly sampled from the set $(-2.0,-0.5)\cup(0.5,2.0)$. The adjacency matrix of the directed edges defines the weights of the structural equation model. Then the data samples are generated using the structural equation, where the noise is sampled from multivariete Gaussian distribution with covariance matrix equals to the adjacency matrix of bidirected edge generated in the previous step.
The third dataset uses real-world data from the financial sector. Paper [1] works with systemic credit risk, one of the most important concerns within the financial system, using dynamic Bayesian networks. The data show that transport and manufacturing companies are likely to transfer risk to other sectors, while banks and the energy sector are likely to be influenced by the risks from other sectors. The data from [1] contains a 10-time series capturing the spreads of 10 European credit default swaps (CDS), and further six time series are added from [45].
Figure 3: Heatmap of weight matrix $W$ (left) and bidirectional weight matrix $W$ (right) on the financial dataset.
We set matrix $F$ to encode for no direct causal relationship between any two pairs of companies from different sectors. Banks sector includes 48DGFE, 05ABBF, 8B69AP, 06DABK, EFAGG9, 2H6677, FH49GG, and 8D8575. Insurance sector includes GG6EBT, DD359M, and FF667M. And lastly, transportation sector and manufacturing includes 0H99B7, 2H66B7, 8A87AG, NN2A8G, and 6A516F.
6.2 Evaluation Criteria
Suppose that a tested algorithm produced weight matrix $\hat{W}$. Such a matrix can contain nearly zero weights. For such reasons, thresholding is done, keeping only edges with weight greater or equal to $\delta$. In cases when the ground-truth weight matrix $W$ is known, the best solution (in terms of structural Hamming distance, see below) is kept over those defined by different threshold $\delta$ values.
In the evaluation, we use the structural Hamming distance (SHD). This distance is the sum of contributions over all pairs of variables in the graph. For two variables $i,j$, let $GT\in\{\rightarrow,\leftarrow,\leftrightarrow,\emptyset\}$ be the edge type in the ground truth graph and $PR\in\{\rightarrow,\leftarrow,\leftrightarrow,\emptyset\}$ be edge type in the predicted graph. Then the contribution of $i,j$ pair to SHD is
$r_{ij}=\begin{cases}0,&\mbox{if }GT=PR,\\ 0.5,&\mbox{if }GT\neq PR\mbox{ and }GT\neq\emptyset\mbox{ and }PR\neq\emptyset,\\ 1,&\mbox{otherwise.}\end{cases}$
Other measured criteria include runtime and F1-score, i.e., the harmonic mean of precision and recall.
6.3 Experiment Setting
In the experiments, we show the results of ExMAG. In the case of synthetic datasets, we generated random graphs with $d\in\{5,10,15,20,25\}$. The number of samples was in $n\in\{20,50,100,500,1000\}$, and ground-truth graph edge ratio was in $\{2,3,4,5,6,7,8,9,10,15,20\}$ percent. Both tested algorithms were run $10$ times, each time with a different seed value and the results were then averaged.
6.4 Experimental Results
The SHD results is shown in Fig. 2 and in the supplementary materials on more datasets. The plots show a comparison of SHD values for ExMAG on both synthetic datasets. As can be seen, the structural Hamming distance grows with the number of variables and decreases with the number of samples. For results with the F1-score, visit the supplementary materials. The results show that the F1 score is close to $1$, with the score increasing with the higher number of samples and decreasing with the higher number of variables.
The results on the real-world dataset can be seen in 3. Contrary to the original expectations, the highest risk importer is company 2H66B7, which stands for Lufthansa. The second highest risk importer is 2H6677, i.e., the Deutsche Bank, which is an expected result.
7 Conclusion
Learning of Dynamic Bayesian networks has received considerable attention as a means of causal learning. With a few exceptions, the research has not considered confounding explicitly. Our method, ExMAG, estimates a maximally ancestral graph, capturing confounding and causal relationships using bidirected and directed edges of a mixed graph. The method provides state-of-the-art statistical performance.
The predictive power of forecasting with variants of dynamical Bayesian networks with confounding considerations is an important direction for further work. While it seems clear that (1) even in dynamical Bayesian networks, marginalization is hard, and thus the computational complexity may be high, (2) the statistical performance should improve by considering the confounding.
Aknowledgements
This work has received funding from the European Union’s Horizon Europe research and innovation programme under grant agreement No. 101084642. The work of G. K. has been partially supported by project MIS 5154714 of the National Recovery and Resilience Plan Greece 2.0 funded by the European Union under the NextGenerationEU Program.
Disclaimer
This paper was prepared for information purposes and is not a product of HSBC Bank Plc. or its affiliates. Neither HSBC Bank Plc. nor any of its affiliates make any explicit or implied representation or warranty and none of them accept any liability in connection with this paper, including, but not limited to, the completeness, accuracy, reliability of information contained herein and the potential legal, compliance, tax or accounting effects thereof. Copyright HSBC Group 2025.
Appendix A F1-score Results
Figure 4: F1-score (in the vertical axis) for different settings of $d$ (in the horizontal axis) and $n$ (horizontal choice of the graph). The plots in the vertical dimension differ according to the dataset used. The channel shows the variance of the results for different choices of seed. Please note that for some of the ER plots, the graphs can be generated only for higher numbers of variables. For example, there exists no ER-5 with $d=10$, as it would need to contain $50$ edges, while maximum is $45$.
Appendix B SHD Results
Figure 5: SHD values (in the vertical axis) for different settings of $d$ (in the horizontal axis) and $n$ (horizontal choice of the graph). The plots in the vertical dimension differ according to the dataset used. The channel shows the variance of the results for different choices of seed.
Appendix C Table of Notation
Table 1: A Table of Notation.
Symbol Representation
$G$ Gender
$D$ Department
$A$ Admit acceptance
$U$ potential confounder, is academic ability in this example
$\mathcal{G}$ Graph
$\mathcal{V}$ Vertex set of the graph
$\mathcal{E}$ Set of directed edges
$\mathcal{U}$ Set of undirected edges
$\text{pa}_{\mathcal{G}}(v)$ Parents of vertex $v$ in graph $\mathcal{G}$
$\text{sp}_{\mathcal{G}}(v)$ Spouses of vertex $v$ in graph $\mathcal{G}$
$\text{an}_{\mathcal{G}}(v)$ Ancestors of vertex $v$ in graph $\mathcal{G}$
$\text{dis}_{\mathcal{G}}(v)$ District of vertex $v$ in graph $\mathcal{G}$
$\text{collider}_{\mathcal{G}}(u)$ Collider nodes in graph $\mathcal{G}$
$W$ Weight matrix
$\hat{W}$ Tested weight matrix
$E$ Directed edge matrix
$B$ Bidirected edge matrix
$F$ Indicating Matrix (existence of a direct causal relationship)
$X_{i,j}$ Value of the $j$-th variable for the $i$-th data point
$w_{k,j}$ Weight of the edge from variable $k$ to variable $j$
$e_{j,k}$ Binary variable indicating a directed edge from $j$ to $k$
$b_{j,k}$ Binary variable indicating a bidirected edge between $j$ and $k$
$f_{j,k}$ Binary variable indicating a existence of directed edge between $j$ and $k$
$d_{j,k}$ Binary variable indicating a directed edge between $j$ and $k$
$r_{ij}$ Contribution of pair $(i,j)$ to SHD
$\lambda$ Regularization parameter
$c$ Large constant for weight bounding
$d$ distance matrix
$n$ Number of data points
$q$ Exponent in the cost function ($q=1$ or $q=2$)
$GT$ Edge type in the ground truth graph
$PR$ Edge type in the predicted graph
$\delta$ Threshold for edge weights
$X,Y$ causal variable
$\theta,\gamma$ regression parameter
$\epsilon_{Y,i},\epsilon_{X,i}$ noise term
$\tilde{w}$ confounding-free regression parameter
$A_{i}$ exogenous variable
$H_{i}$ unobserved term
$\mathcal{P}$ class of distributions
$C$ conditioning set