### v3

# ExMAG: Learning of Maximally Ancestral Graphs

Abstract
In mixed graphs, there are both directed and undirected edges. An extension of acyclicity to this mixed-graph setting is known as maximally ancestral graphs. This extension is of considerable interest in causal learning in the presence of confounders. There, directed edges represent a clear direction of causality, while undirected edges represent confounding.
We propose a score-based branch-and-cut algorithm for learning maximally ancestral graphs. The algorithm produces more accurate results than state-of-the-art methods, while being faster to run on small and medium-sized synthetic instances.
1 Introduction
As one transitions from statistical to causal learning [50], one is seeking the most appropriate causal model. Dynamic Bayesian networks (DBN) [19, 38] are a popular model, where a weighted directed acyclic graph represents the causal relationships. Stochastic processes are represented by their vertices, and weighted, oriented edges suggest the strength of the causal relationships. The key challenge in learning DBNs is confounding.
To illustrate the challenge of confounding, let us consider Simpson’s paradox. Simpson’s paradox shows that without considering confounding factors in statistical analysis [36], the direction of causality can be mis-estimated completely. An textbook example [36] comes from the Berkeley graduate admissions [4]. The data show that women find it harder to get admitted to Berkeley graduate schools. Nevertheless, this is because women tend to apply to departments that have lower admission rates. In this example, the choice of the graduate school is the confounder, impacting the probability of admission. Confounding is prevalent throughout high-dimensional statistics [34, 23], such as in biomedical sciences.
Specifically, in biomedical sciences, confounders such as socio-economic status, age, or lifestyle factors can distort the true causal relationship between treatments and outcomes [55]. Techniques such as instrumental variables [42, 31], propensity score matching [46], and double machine learning [12] have been widely used to mitigate the effects of confounding in clinical trials and observational studies. In bioinformatics, particularly in genome-wide association studies (GWAS), confounders, including population stratification and environmental exposures, must be controlled to avoid biased estimates in genetic association studies [10, 24]. To mitigate such biases, statistical models that explicitly account for hidden confounders, such as spectral methods and latent variable models, are often employed [24]. Furthermore, meta-analysis and sensitivity analysis are often employed to evaluate the robustness of findings in the presence of potential confounders, especially when combining results from multiple studies [7, 35]. These methodologies ensure that the conclusions drawn are reliable and actionable, improving the credibility of statistical models across disciplines.
In statistical theory, work [9] studies confounding in detail, and many subsequent works develop further methods. [33] shows that leveraging the dominant eigenstructure of time series may improve performance of estimation. Anchor regression, for instance, bridges the gap between causality and robustness by addressing heterogeneity in data [47]. Other significant contributions include spectral deconfounding, a technique designed to mitigate the effects of hidden confounders in high-dimensional settings [8]. This approach provides a framework for robust predictions in the presence of shifts in data distributions. Similarly, the invariance principle has emerged as a cornerstone of causal inference, linking causal structure to robust statistical models [7]. Furthermore, the concept of doubly robust inference offers an alternative framework for addressing hidden confounding factors, combining model robustness with efficiency in high-dimensional scenarios [24]. Together, these developments represent a significant step forward in understanding and addressing the challenges posed by complex causal systems.
While there is a long history of the study of confounding, as suggested above, the extensions of DBNs to allow for confounding are rather more recent. Instead of estimating a Directed Acyclic Graph (DAG), one could estimate a Maximal Ancestral Graph (MAG), cf. [44]. MAGs allow for both direct and indirect relationships among variables modelled as directed and undirected edges, even in the presence of confounding factors. In particular, MAGs can represent feedback loops and bidirectional relationships that DAG-based models, such as DBNs, cannot. This makes MAGs a more powerful tool for capturing the complex dynamics of real-world causal systems.
There are only a few studies of MAG estimation [11, 41, 13, 26, 29, 28, 17]. [43, 53] are applicable to both discrete and nonparametric cases, which extend DAG to MAG or ADMG diagrams, see more definitions in Sect. 2. Factorization in MAG is not directly decomposable into individual variables and their parent sets, as in DAGs, but must instead consider components connected by bidirected paths (termed districts or c-components), cf. [43], although [13] proposed to use Markov equivalence classes (MEC) instead. In 2021, [11] introduced a first mixed-integer programming (MIP) formulation, but the number of variables scales with the number of c-components, i. e., exponentially with the number of vertices in the worst case. Such formulations are also known as extended formulations [14]. The same year, [41] explored a constraint-based approach for MAG discovery, leveraging conditional independence testing. Additionally, [56] addresses exogenous covariates in causal formulation that helps explain the heterogeneity in both sampling and causal mechanisms. Dissertation [26] presented an extension of the imsets of [52] from directed acyclic graphs (DAGs) to towards MAGs [29], which allows for the use of the methods of Studený, and a score-based heuristic [28]. More recently, paper [17] enhanced the scalability of methods of [11] by utilizing linear programming (LP) relaxations instead of solving the MIP.
Our approach proposes a formulation of MAG estimation within Mixed-Integer Nonlinear Programming (MINLP) in a dimension polynomial in the number of vertices, in contrast with the so-called extended formulations of [11, 17], where the dimension is exponential in the number of vertices. While both the extended formulation of [11] and ours ensure that confounding factors are properly accounted for and the true underlying data-generating process is better represented by the model, our implementation scales further, from 4-5 stochastic processes in the extended formulation of [11] to 25 or more stochastic processes with the proposed compact formulation.
AbilityDepartmentAdmitGender
(a) Cofounding graph
[…]
(b) Result of ExDBN
[…]
(c) Result of ExMAG
Figure 1: Ground truth with the confounder of Department on the Berkeley graduate admission example (left, 1(a)), a dynamic Bayesian network trained on the data (center, 1(b)), and ExMAG output (right, 1(c)). While the dynamic Bayesian network suggests a causal relationship between gender and admission, ExMAG correctly identifies the confounding. See the supplementary material for details.
1.1 Motivating Example
Let us revisit the Berkeley graduate admission paradox example from the first page. As in most paradoxes, there is no violation of logic in Simpson’s paradox, just a violation of of intuition. The poor intuition being violated in this case is that a positive association in the entire population should also hold within each department. Overall, females in these data did have a harder time getting admitted to graduate school. But that arose, because female applicants chose the departments that were the most difficult to gain admission to for anyone, male or female. In this example, gender influences the choice of department, and the department influences the chance of admission. Controlling for department reveals a more plausible direct causal influence of gender, as illustrated in the DAG in Fig. 1(a). Our method, ExMAG, is able to reveal the confounders in this Berkeley graduate admission example, as illustrated in the notebook in supplementary materials and Figure 1.
2 Graphs and Properties
A DAG is a directed graph $G=(V,E)$ such that there are no directed cycles. That is, there is no sequence of distinct vertices $v_{1},v_{2},\dots,v_{k}\in V$ such that $(v_{i},v_{i+1})\in E$ for all $11\leq i\leq k-$ 1 and $(v_{k},v_{1})\in E$. Maximal ancestral graphs (MAGs), first introduced by [44], provide a framework for modeling distributions through conditional independence (CI) relations. Compared with directed acyclic graphs (DAGs), MAGs allow for latent confounders, accommodating data that arise from distributions with more complex independence structures and revealing hidden states in the graphs. While DAGs allow for the efficient computation of maximum likelihood estimates (MLEs) and scoring (e.g., via BIC), these properties are challenging to extend to MAG due to their structural and computational complexity [27].
ADMG
Mixed graphs feature two types of edges: directed ($\rightarrow$) and bidirected ($\leftrightarrow$). Mixed graph $\mathcal{G}$ thus consists of a vertex set $\mathcal{V}$, a set of directed edges $\mathcal{E}$ and undirected edges $\mathcal{U}$, where $\mathcal{E}$ are (ordered) pairs of distinct vertices, while $\mathcal{U}$ are (unordered) 2-element subsets of vertices. For a directed edge in $\mathcal{E}$ connecting vertices $v$ and $w$, we say these two vertices are the endpoints of the edge and the two vertices are adjacent (if there is no edge between $v$ and $w$, they are non-adjacent). For a vertex $v$ in $\mathcal{V}$, we define the parents, spouses, ancestors, and district of $v$, respectively as:
$\text{pa}_{\mathcal{G}}(v)$ =$=\{w:w\to v\text{ in }\mathcal{G}\},$ $\text{an}_{\mathcal{G}}(v)$ =$=\{w:w\to\cdots\to v\text{ in }\mathcal{G}\text{ or }w=v\},$
$\text{sp}_{\mathcal{G}}(v)$ =$=\{w:w\leftrightarrow v\text{ in }\mathcal{G}\},$ $\text{dis}_{\mathcal{G}}(v)$ =$=\{w:w\leftrightarrow\cdots\leftrightarrow v\text{ in }\mathcal{G}\text{ or }w=v\}.$
Given a directed mixed graph $\mathcal{G}$, the districts define a set of equivalence classes of nodes in $\mathcal{G}$. The district for node $v$ is defined as the connected component of $v$ in the subgraph of $\mathcal{G}$ induced by all bidirected edges. As in a DAG, a mixed graph $\mathcal{G}$ is acyclic if it contains no directed cycles in $\mathcal{E}$, i.e., an acyclic directed mixed graph (ADMG) [26].
A directed mixed graph G is called an ancestral ADMG if the following condition holds for all pairs of nodes $v$ and $w$ in $\mathcal{G}$:
$\text{If }v\neq w\text{ and }v\in\text{an}_{\mathcal{G}}(w)\cup\text{sp}_{\mathcal{G}}(w),\text{ then }w\notin\text{an}_{\mathcal{G}}(v),$
which is written as, $\mathcal{G}$ is an ancestral ADMG if it contains no directed cycles ($v\to u\to\ldots\to w\to v$) or almost directed cycles [11, 26]. In an ADMG, an almost directed cycle is of the form $v\to u\to\ldots\to w\leftrightarrow v$; in other words, $\{v,w\}\in\mathcal{U}$ is a bidirected edge, and $v\in\text{an}_{\mathcal{G}}(w)$ [11].
Inducing Paths
An inducing path from variable $X$ to variable $Y$ in a directed graph $G=(V,E)$ is a path $P=(X=v_{0},v_{1},\dots,v_{n}=Y)$ such that for all intermediate nodes $v_{i}$ (where $11\leq i\leq n-$ 1), $v_{i}\in Z$, where $Z$ is the conditioning set. If the path is blocked by conditioning on $Z$, then it is an inducing path.
A node $u$ on a non-overlapping path is called a collider if contains a non-overlapping subpath $(w,u,v)$ with two arrowheads into $u$. In mathematical form, a collider is represented as
$\text{collider}_{G}(u)=\left\{u:\exists w,v:\begin{array}[]{c}w\to u\leftarrow v\vee w\leftrightarrow u\leftarrow v\vee w\to u\leftrightarrow v\vee w\leftrightarrow u\leftrightarrow v\end{array}\right\}.$
m-separation
Graphs encode conditional independence via separation criteria. For acyclic directed mixed graphs (ADMGs), m-separation generalizes d-separation to handle bidirected edges. A path between nodes $a$ and $b$ is m-connecting given a conditioning set $C\subseteq\mathcal{V}$ if: (i) $a$ and $b$ are the endpoints; (ii) all non-colliders are not in $C$; and (iii) all colliders are in $\operatorname{an}_{G}(C)$. Nodes $a$ and $b$ are m-separated given $C$ if no such path exists.
Maximal Ancestral Graph
An ADMG $\mathcal{G}$ is called a maximal ancestral graph (MAG) if:
  1. (i)
    For every pair of nonadjacent vertices $a$ and $b$, there exists some set $C$ such that $a,b$ are m-separated given $C$ in $\mathcal{G}$ (Maximality);
  2. (ii)
    For every $v\in\mathcal{V}$, $\text{sib}_{\mathcal{G}}(v)\cap\text{anc}_{\mathcal{G}}(v)=\emptyset$ (Ancestrality).
where $\text{sib}_{\mathcal{G}}(v)=\{u\in\mathcal{V}\mid\exists w\in\mathcal{V}:w\to v,\;w\to u,\;u\not\to v,\;v\not\to u\}$. We refer to [26] for multiple examples.
3 Formulation of the Mixed Integer Quadratic Program
Recent works on high-dimensional confounding or deconfounding clarify the connections between distributional robustness, replicability, and causal inference [47, 24]. Distributional robustness differs significantly from traditional robust statistical methods [30, 25], which typically handle outliers in the training data, while our work focuses on evaluating the existence of a confounding factor.
In this section, we inherit from distributional robustness and present the formulation of the Mixed Integer Quadratic Program (MIQP) to infer the causal structure. Since we cannot observe all relevant variables, we must deal with the situation of hidden confounding. The problem is formalized in the following form corresponding to a structural equation model (SEM) [5, 39]:
$Y_{i}\leftarrow X_{i}\tilde{w}+g(H_{i},A_{i})+\epsilon_{Y,i},$
where:
  • $i$ indexes the sample instances;
  • $\epsilon_{Y,i}$ is the noise term, independent of all variables that appear ”upstream” from $Y_{i}$.
  • $A_{i}$ is an exogenous variable, though not considered in the following sections.
  • $\leftarrow$ represents an algebraic equality, implying a structural causal relationship.
  • $H_{i}$ is the unobserved confounding variable vector.
If non-zero components of the vector $\tilde{w}$ are correlated with certain components of $X_{i}$, these components are defined as causal $X$-variables for $Y_{i}$. This means:
[…]
For the scenario that there are no exogenous variable perturbations, we describe this with an additive noise model with hidden states as follows:
$Y_{i}$ $\leftarrow X_{i}\tilde{w}+H_{i}\theta+\epsilon_{Y,i},$ (2)
$X_{i}$ $\leftarrow H_{i}\gamma+\epsilon_{X,i},$ (3)
where $\epsilon_{X,i}$, $\epsilon_{Y,i}$ and $H_{i}$ are mutually independent. $X_{i}$ is the observed $p\times 1$ covariate vector. $H_{i}$ is the unobserved $q\times 1$ confounding variable vector. Our goal is to infer the confounding-free regression parameter $\tilde{w}$ and stabilize the prediction of the relationship between $Y$ and $X$.
3.1 Connecting to Causality
The causal parameter $\tilde{w}$ can be seen as minimizing the worst-case risk:
$\tilde{w}=\arg\min_{w}\max_{P\in\mathcal{P}}\mathbb{E}_{P}\left[(Y_{i}-X_{i}w)^{2}\right],$
where $\mathcal{P}$ is a class of distributions containing perturbations of the original distribution, including confoundings. This modeling highlights the inherent connection between causality and distributional robustness [18, 40, 45].
In this perspective, we present the formulation of the Mixed Integer Quadratic Program (MIQP) used to infer the causal structure, with a new binary matrix $B=[b_{j,k}]\in\{0,1\}^{d\times d}$ introduced to account for relationships explained by confounding factors, alongside the weight matrix $W$ and binary adjacency matrix $E$ adopted from the ExDAG model [49]. Weight matrix $W$ represents the model weights. Whenever an entry in $W_{j,k}$ is non-zero, the respective value in either $E_{j,k}$ (directed edge) or $B_{j,k}$ is nonzero (bidirected edge). At the same time, we extend the existing formulation by introducing an additional binary input parameter $f_{j,k}$ for each pair of variables $(j,k)$, where $j\neq k$, indicating that there is no direct causal relationship between variables $j$ and $k$, but $j$ and $k$ might have a common cofactor. This follows from the meaning of the edges in a MAG $\rightarrow$ edge implies a direct causal relationship but does not rule out a possible latent confounding, $\leftrightarrow$ means no direct causal relationship.
Formally, we define Directed Edge Matrix $E$ as $e_{j,k}=1$ if $j\to k$, 00 otherwise. Similarly, Bidirected Edge Matrix $B$ is $b_{j,k}=1$, and 00 otherwise. Lastly, input matrix $F$ with the set of variables known to be independent, is by definition $f_{j,k}=1$ if $j\not\to k\wedge k\not\to j$ and 00 otherwise. Based on these assumptions, one can substitute (3) into (2), and the work of [49] is extended as explained in the next section.
3.2 MILP Formulation
The cost function for the Mixed Integer Quadratic Program of ExMAG is the following $l_{q}$ norm:
$l_{q}$ $=\sum_{i=1}^{n}\sum_{j=1}^{d}\left[X_{i,j}-\sum_{k=0;k\neq j}^{d}X_{i,k}w_{k,j}\right]^{q}+\lambda\sum_{j=0}^{d}\sum_{k=0}^{d}(e_{j,k}+b_{j,k}),$ (4)
where:
  • $X_{i,j}$ represents the value of the $j$-th variable for the $i$-th data point;
  • $w_{k,j}$ represents the weight of the edge from variable $k$ to variable $j$;
  • $e_{j,k}$ is the binary decision variable indicating the presence of a directed edge from $j$ to $k$;
  • $b_{j,k}$ is the binary decision variable indicating a bidirected edge between $j$ and $k$;
  • $\lambda$ is a regularization parameter controlling the model fit and the edge penalty trade-off.
  • The exponent $q\in\mathbb{N}$ can take values $q=1$ or $q=2$.
As in the ExDAG [48] model, the weights are bounded by introducing a large constant $c$.The bounding avoids bilinear terms in the cost function in (4) and takes the following form:
$-c\cdot(E+B)$ $\leq W\leq c\cdot(E+B)$ (Weight Constraint)
$E+B$ $\leq\mathbf{1}$
The 3.2 means that there cannot be a directed as well as a bidirected edge between the same two vertices. Additionally, we enforce that the bidirected matrix is symmetric by (5). If $f_{j,k}=1$, then we know there is no direct causal relationship between $j$ and $k$, and therefore, $e_{j,k}=0$. This is formally enforced by (6) Inversely, $f_{j,k}=0$ implies a directed edge rather than a bidirected edge between $j$ and $k$ in (7).
$B$ $=B^{T},$ (5)
$F+E$ $\leq\mathbf{1},$ (6)
$B$ $\leq F.$ (7)
Lastly, we must enforce conditions for directed or almost directed cycles and inducing paths. Those conditions are enforced lazily using a separation routine explained later. Directed cycles are enforced in a way adopted from [49]. Therefore, they are left out of this paper. An almost directed cycle formed by edges in set $E^{\prime}$ and a bidirected edge $(u,v)$ is forbidden by the constraint
$b_{u,v}+\sum_{(j,k)\in E^{\prime}}e_{j,k}\leq|E^{\prime}|.$
Similarly, if there is an inducing path formed by path $P$ that contains bidirected edges, and set $E^{\prime}$ contains all directed edges that participate in the ancestor relationship (including multiple paths) between the inner points of the path and the terminals of $P$, this inducing path is forbidden by
$\sum_{(j,k)\in P}b_{j,k}+\sum_{(j,k)\in E^{\prime}}e_{j,k}\leq|E^{\prime}|+|P|-1.$
Note that the second condition does not necessarily eliminate the inducing path, as the optimizer might forbid one of the edges in $E^{\prime}$ without influencing the ancestor relationship. This results in path $P$ being found in the next iteration, with a smaller set of directed edges, and the process is repeated.
By enforcing these constraints, we ensure that the MIQP correctly models the causal relationships between the variables while respecting the independence structure defined by $f_{j,k}$ and the potential confounding relationships captured by $b_{j,k}$.
4 Separation Routine for the Maximal Ancestral Graphs
The main contribution of this section is the separation routine that identifies whenever a graph is an instance of a maximal ancestral graph. To do so, we need to identify directed cycles, almost directed cycles, and inducing paths. The presence of directed cycles can be detected in $\mathcal{O}(d^{2})$ using depth-first-search (DFS); such an approach can be found in [49]. For both inducing paths and almost directed cycles, we will use the distance matrix $D$ constructed on the graph of directed edges $E$. This distance matrix can be obtained, for example, using the Floyd-Warshall algorithm [22].
Having the distance matrix, to check for almost directed cycles, we can iterate over all bidirected edges and test whether the distance between the endpoints using $E$ is finite, i.e., we have a directed path connected by a bidirected edge. See Algorithm 1 for details.
Input: directed edges $E$, bidirected edges $B$
function Almost-Directed-Cycles($E$, $B$)
     $D\leftarrow$ Distance-Matrix($E$)
     for all $(j,k)\in\{1,2,\ldots,d\}\times\{1,2,\ldots,d\}$ do
         if $j\neq k\And b_{j,k}==1\And D_{j,k}<\infty$ then
              $E^{\prime}=$ Trace-Distance-Matrix($D$,$E$,$j$,$k$) $\triangleright$ Finds all edges on any $j$ to $k$ path, see Supl.
              Found cycle formed by edges $E^{\prime}$ and $j\leftrightarrow k$
         end if
     end for
end function
Algorithm 1 Function that identifies almost directed cycles.
In the case of inducing paths, we use a DFS starting from each vertex. Once started from vertex $s$, the DFS routine checks for all possible inducing paths that terminate in $s$. For efficiency, a set of all possible endpoints of the path is held. Once this set is empty, the DFS search is terminated, and no further exploration is performed. The set is updated using the distance matrix calculated on the directed edges. If we consider a vertex $v$, $s$ must be either its ancestor (meaning that possible endpoints for $v$ remain unchanged) or the second inducing path endpoint is among the points that are reachable from $v$ (meaning that the possible endpoints for $v$ are replaced with their intersection with the set of all points reachable from $v$). See Algorithm 2 for details.
Input: directed edges $E$, bidirected edges $B$
function Inducing-Paths($E$, $B$)
     $D\leftarrow$ Distance-Matrix($E$)
     for all $s\in{1,2,\ldots,d}$ do Inducing-Paths-DFS($D$, $E$, $B$, $s$, $s$, $\{1,2,\ldots d\}$, $[s]$)
     end for
end function
function Inducing-Paths-DFS($D$, $E$, $B$, $s$, $u$, possible endpoints, path)
     if possible endpoints are empty then return
     end if
     if Len(path) $>2\And u$ in possible endpoints  then Found-Inducing-Path($D$,$E$, path)
         $\triangleright$ Edges participating in ancestor relation are recovered, see Suppl. mat.
     end if
     for all $v\in{1,2,\ldots,d}$ such that $e_{u,v}=1$ do
         v-endpoints $\leftarrow$ possible endpoints if $D_{v,s}<\infty$ else possible endpoints $\cap\{x\mid D_{v,x}<\infty\}$
         Inducing-Paths-DFS($D$,$E$,$B$,$s$,$v$,v-endpoints,path$+v$)
     end for
end function
Algorithm 2 Function that identifies inducing paths.
Once having the bidirected edges in the inducing paths and almost directed cycles, we need to trace back the Floyd-Warshall distance matrix to find all directed edges that form the cycle or the ancestor relationship. This is done using calls to the function Trace-Distance-Matrix, which can be found in the Supplementary materials.
If directed cycles, almost directed cycles, and inducing paths are found, the algorithm applies lazy constraints in 3.2 and in 3.2. Note that removing one directed edge between two vertices where multiple paths exist is not a necessary condition for the graph to become MAG; however, this procedure can be repeated iteratively. If no inducing paths or almost directed cycles are found, we know that the program converged to the optimum, and we have a maximal ancestral graph, which minimizes (4).
5 Experimental Evaluation
Figure 2: SHD values (in the vertical axis) for different settings of $d$ (in the horizontal axis) and $n$ (horizontal choice of the graph). The plots in the vertical dimension differ according to the dataset used. Standard deviations are depicted as the blurred regions, and dashed lines are the maximum values. See supplementary materials for results on more datasets and error information.
Figure 3: SHD values on 3BF datasets.
Figure 4: Run times of the compared algorithms in seconds.
Used Datasets
We tested the ExMAG algorithm on both synthetic and real-world datasets. The first synthetic data set is based on the Erdös-Rényi model (ER) [21], in which the ground truth graph is randomly selected from all graphs with $d$ vertices and $m$ edges (parameter of the experiment, for example, dataset ER-2 contains $2$ edges per variable, that is, $m=2\cdot d$). The weights of the graph are randomly sampled from the set $(-2.0,-0.5)\cup(0.5,2.0)$.
Once the ground truth model is created, the training data are generated using the structural model equation (2),  (3) ($H_{i}$ set to 00). Then, $20\,\%$ of variables are treated as latent variables and hidden from the training data. The respective columns and rows from the ground truth weight matrix $W$ have also been removed. Finally, $20\,\%$ of edges between variables not connected by an edge in the ground truth data are marked in $F$.
The second dataset uses randomly generated bow-free (BF) graphs, a subset of all possible MAGs. A bow-free graph is a graph such that for no pair of vertices $i,j$, $i\rightarrow j$ and at the same time $i\leftrightarrow j$. The BF graph generation process has two parameters: the probability of a directed edge and the probability of a bidirected edge. The generation process is as follows. First, a bow-free graph with the given edge probabilities is generated randomly. Then the weights of the sampled graphs are randomly sampled from the set $(-2.0,-0.5)\cup(0.5,2.0)$.
The third synthetic dataset, 3BF, is a modified version of BF. We generate the ground truth graph in the same way as in the BF dataset and then modify it. Specifically, we identify each vertex with a degree greater than three, and randomly remove edges until the vertex has a degree of at most three.
The adjacency matrix of the directed edges defines the weights of the structural equation model. Then the data samples are generated using the structural equation, where the noise is sampled from a multivariate Gaussian distribution with a covariance matrix equal to the adjacency matrix of bidirected edges generated in the previous step.
The fourth dataset uses real-world data from the financial sector. Paper [1] works with systemic credit risk, one of the most important concerns within the financial system, using dynamic Bayesian networks. The data show that transport and manufacturing companies are likely to transfer risk to other sectors, while banks and the energy sector are likely to be influenced by the risks from other sectors. The data from [1] contains a 10-time series capturing the spreads of 10 European credit default swaps (CDS), and further six time series are added from [48].
We set matrix $F$ to encode for no direct causal relationship between any two pairs of companies from different sectors. Banks sector includes 48DGFE, 05ABBF, 8B69AP, 06DABK, EFAGG9, 2H6677, FH49GG, and 8D8575. Insurance sector includes GG6EBT, DD359M, and FF667M. And lastly, transportation sector and manufacturing includes 0H99B7, 2H66B7, 8A87AG, NN2A8G, and 6A516F.
Evaluation Criteria
Suppose that a tested algorithm produced weight matrix $\hat{W}$. Such a matrix can contain nearly zero weights. For such reasons, thresholding is done, keeping only edges with weight greater than or equal to $\delta$. In cases when the ground-truth weight matrix $W$ is known, the best solution (in terms of structural Hamming distance, see below) is kept over those defined by different threshold $\delta$ values. In the evaluation, we use the structural Hamming distance (SHD). This distance is the sum of contributions over all pairs of variables in the graph. For two variables $i,j$, let $GT\in\{\rightarrow,\leftarrow,\leftrightarrow,\emptyset\}$ be the edge type in the ground truth graph and $PR\in\{\rightarrow,\leftarrow,\leftrightarrow,\emptyset\}$ be edge type in the predicted graph. Then the contribution of $i,j$ pair to SHD is $r_{ij}=0$ if $GT=PR$, $0.5$ if $GT\neq PR\wedge GT\neq\emptyset\wedge PR\neq\emptyset$, and $1$ otherwise. Other measured criteria include runtime and F1-score, i.e., the harmonic mean of precision and recall.
Experiment Setting
In the experiments, we show the results of ExMAG. In the case of synthetic datasets, we generated random graphs with the number of vertices $d\in\{5,10,15,20,25\}$. The number of samples was $n\in\{20,50,100,500,1000\}$, and for the ground-truth graph, the edge-to-vertex ratio was in $\{2,3,4,5,6,7,8,9,10,15,20\}$. All tested algorithms were run $10$ times, each time on synthetic data generated using a random generator initialized with a different seed. The results were then averaged. We compared our method with the FCI algorithm [51], ip4ancadmg [11], and ipbm [17]. We set regularization coefficient $\lambda$ to $1.0$. We ran experiments on a computing cluster with AMD EPYC 7543 cpus and each job had allocated two cores and 64GB RAM. Time limit was 900 seconds for ExMAG and 1800 seconds for other methods. The total cpu time needed for experiments in this paper was around one month.
Figure 5: Heatmap of weight matrix $W$ (left) and bidirectional weight matrix $W$ (right) on the financial dataset.
Experimental Results
The SHD results are shown in Figures 2 and 3 and in the supplementary materials for additional datasets. The plots show a comparison of SHD values for ExMAG on the synthetic datasets. As can be seen, the structural Hamming distance grows with the number of variables and decreases with the number of samples.
As we can see on Figure 2 and 3, ExMAG performs better than FCI on all scenarios. Since both ipbm and ip4ancadmg have a preprocessing step that depends exponentially on the maximum in-degree of the underlying ground truth graph, we tested these two algorithms only on the 3BF datasets, where the in-degree is bounded by three. We can see in Figure 3, that ExMAG also performs better than ipbm and ip4ancadmg. The run times of the evaluated algorithms are shown in Figure 4. For additional results (incl. the F1-score), please see the supplementary materials.
The results on the real-world dataset can be seen in Figure 5. Contrary to the original expectations, the highest risk importer is company 2H66B7, which stands for Lufthansa. The second highest risk importer is 2H6677, i.e., the Deutsche Bank, which is an expected result.
6 Conclusion and Limitations
Learning of dynamic Bayesian networks has received considerable attention as a means of causal learning. With a few exceptions, the research has not considered confounding explicitly. Our method, ExMAG, estimates a maximally ancestral graph, capturing confounding and causal relationships using bidirected and directed edges in a mixed graph. The method provides state-of-the-art statistical performance.
As with many other methods for causal learning, the scalability of the method may leave space for improvement. Although the branch-and-cut algorithm runs in time that is exponential in the number of time series in the worst case, Figure 4 illustrates that our run time is lower than those of ip4ancadmg [11] and ipbm [17], while improving the SHD of both recent competitors at the same time (cf. Figure 3). One could improve upon the run-time further by introducing additional cutting planes and more elaborate data structures for the separation of 3.2 and 3.2, perhaps drawing inspiration from solvers [15, e.g.] for the travelling salesman problem.
In terms of future work, exploring the predictive power of forecasting using variants of dynamical Bayesian networks with confounding considerations seems prominent. Although it seems clear that marginalization is hard even in dynamical Bayesian networks, and thus the computational complexity may be high, but statistical performance is likely to improve, when confounding is considered.
Aknowledgements
This work has received funding from the European Union’s Horizon Europe research and innovation programme under grant agreement No. 101084642. The work of G. K. has been partially supported by project MIS 5154714 of the National Recovery and Resilience Plan Greece 2.0 funded by the European Union under the NextGenerationEU Program.
Disclaimer
This paper was prepared for information purposes and is not a product of HSBC Bank Plc. or its affiliates. Neither HSBC Bank Plc. nor any of its affiliates make any explicit or implied representation or warranty and none of them accept any liability in connection with this paper, including, but not limited to, the completeness, accuracy, reliability of information contained herein and the potential legal, compliance, tax or accounting effects thereof. Copyright HSBC Group 2025.
Appendix A F1-score Results
Figure 6: F1-score (in the vertical axis) for different settings of $d$ (in the horizontal axis) and $n$ (horizontal choice of the graph). The plots in the vertical dimension differ according to the dataset used. Standard deviations are depicted as the blured regions and dashed lines are the minimum values. Please note that for some of the ER plots, the graphs can be generated only for higher numbers of variables. For example, there exists no ER-5 with $d=10$, as it would need to contain $50$ edges, while maximum is $45$.
Figure 7: F1-score comparison of all evaluated algorithms on 3BF datasets.
Appendix B SHD Results
Figure 8: SHD values (in the vertical axis) for different settings of $d$ (in the horizontal axis) and $n$ (horizontal choice of the graph). The plots in the vertical dimension differ according to the dataset used. Standard deviations are depicted as the blured regions and dashed lines are the maximum values.
Appendix C Table of Notation
Table 1: A Table of Notation.
Symbol Representation
$G$ Gender
$D$ Department
$A$ Admit acceptance
$U$ potential confounder, is academic ability in this example
$\mathcal{G}$ Graph
$\mathcal{V}$ Vertex set of the graph
$\mathcal{E}$ Set of directed edges
$\mathcal{U}$ Set of undirected edges
$\text{pa}_{\mathcal{G}}(v)$ Parents of vertex $v$ in graph $\mathcal{G}$
$\text{sp}_{\mathcal{G}}(v)$ Spouses of vertex $v$ in graph $\mathcal{G}$
$\text{an}_{\mathcal{G}}(v)$ Ancestors of vertex $v$ in graph $\mathcal{G}$
$\text{dis}_{\mathcal{G}}(v)$ District of vertex $v$ in graph $\mathcal{G}$
$\text{collider}_{\mathcal{G}}(u)$ Collider nodes in graph $\mathcal{G}$
$W$ Weight matrix
$\hat{W}$ Tested weight matrix
$E$ Directed edge matrix
$B$ Bidirected edge matrix
$F$ Indicating Matrix (existence of a direct causal relationship)
$X_{i,j}$ Value of the $j$-th variable for the $i$-th data point
$w_{k,j}$ Weight of the edge from variable $k$ to variable $j$
$e_{j,k}$ Binary variable indicating a directed edge from $j$ to $k$
$b_{j,k}$ Binary variable indicating a bidirected edge between $j$ and $k$
$f_{j,k}$ Binary variable indicating a existence of directed edge between $j$ and $k$
$d_{j,k}$ Binary variable indicating a directed edge between $j$ and $k$
$r_{ij}$ Contribution of pair $(i,j)$ to SHD
$\lambda$ Regularization parameter
$c$ Large constant for weight bounding
$d$ Number of variables
$n$ Number of data points
$q$ Exponent in the cost function ($q=1$ or $q=2$)
$GT$ Edge type in the ground truth graph
$PR$ Edge type in the predicted graph
$\delta$ Threshold for edge weights
$X,Y$ causal variable
$\theta,\gamma$ regression parameter
$\epsilon_{Y,i},\epsilon_{X,i}$ noise term
$\tilde{w}$ confounding-free regression parameter
$A_{i}$ exogenous variable
$H_{i}$ unobserved term
$\mathcal{P}$ class of distributions
$C$ conditioning set
Appendix D Pseudocode
Input: distances $D$ defined by directed edges $E$, start point $j$, and endpoint $k$
Output: edges on any path from $j$ to $k$
function Trace-Distance-Matrix($D$, $E$, $j$, $k$)
     if $D_{j,k}==\infty$ then return
     end if
     visited = $\{(j,k)\}$
     stack = stack with $(j,k)$
[…]
     while stack is not empty do
         $u,v\leftarrow$ Pop(stack)
         for all $w\in{1,2,\ldots,d}$ s t. $D_{u,w}+D_{w,v}<\infty$ do
              visited $\leftarrow$ visited $\cup\{(u,w),(w,v)\}$
              if $E_{u,w}$ then edges $\leftarrow$ edges $\cup\{(u,w)\}$
              else if $E_{w,v}$ then edges $\leftarrow$ edges $\cup\{(w,v)\}$
              end if
              add $\{(u,w),(w,v)\}$ to stack
         end for
     end whilereturn edges
end function
function Found-Inducing-Path($D$,$E$, $P$)
[…]
     for all vertices $j\in$ $P$ and $j\not\in\{P_{0},P_{|P|}\}$ do
         $E^{\prime}=E^{\prime}\cup$ Trace-Distance-Matrix($D$,$E$,$j$,$P_{0}$) $\cup$ Trace-Distance-Matrix($D$,$E$,$j$,$P_{|P|}$)
         $\triangleright$ Finds all edges on any $j$ to $P_{0}$ ($P_{|P|}$) path
     end for
     Found inducing path formed by path $P$ and directed edges $E^{\prime}$
end function
Algorithm 3 Functions that help in the separation routine.
Appendix E Residual Assumptions And Statistical Significance
There exists a fundamental premise in structural equation modeling that the residuals—representing unexplained variation—are asymptotically unbiased, meaning they are independent of both observed and latent variables, and follow a zero-mean distribution. This assumption plays a critical role in ensuring that learned causal relationships are not distorted by hidden confounders or systematic error. The ExMAG framework embraces this principle by design, explicitly modeling residual independence as a safeguard against spurious causal edges. Just as fairness-aware systems aim to isolate structural patterns from social bias [2], ExMAG works to separate signal from statistical noise. The result is a model capable of learning causal mechanisms that are not only mathematically sound but also resilient across different subpopulations, forming a foundation of causal inference.
Causal discovery systems, like decision-making algorithms in high-stakes domains, must operate effectively across structurally diverse populations. This paper uses real-world financial data spanning multiple sectors—banking, insurance, manufacturing, and transportation—each exhibiting distinct systemic exposures. These domains can be viewed as a privileged setting where data availability and quality are high, yet subgroup heterogeneity remains significant. In such contexts, ExMAG successfully identifies dominant risk propagation patterns, even when feature distributions vary across industries. This mirrors broader challenges in fairness: the need to perform robustly across populations with unequal baseline conditions [37]. The model’s consistent recovery of risk links—illustrated in Figure 5—not only affirms its structural fidelity but also its capacity to generalize without group-specific tuning.
Understanding the statistical reliability of a model’s output requires more than average performance—it demands insight into variance. To that end, the authors conduct 10 independent trials for each configuration, reporting both mean and standard deviation for key metrics such as SHD and F1-score. The inclusion of error bars in Figures 2 and 6 provides a visual representation of variability, revealing not just how well the model performs, but how consistently. In contrast to baseline methods with large fluctuations, ExMAG demonstrates narrow error margins, underscoring its stability in the face of stochastic elements like data partitioning and initialization.
Appendix F Brief Introduction To Mixed Integer Quadratic Programming
Let us also provide a short introduction to mixed-integer quadratic programming. An optimization problem is called a mixed-integer quadratically constrained quadratic program (MIQCQP) if it is of the form
$\underset{x\in\mathbb{R}^{n}}{\min}\quad$ $x^{T}Qx+q^{T}x,$ (8)
s.t. $x^{T}Q_{i}x+q_{i}^{T}x\leq a_{i},$ (9)
$Ax\leq b,$ (10)
$x\in\mathbb{R}^{n-r}\times\mathbb{Z}^{r}$ (11)
where $Q,Q_{i}\in\mathbb{R}^{n\times n}$, $q,q_{i}\in\mathbb{R}^{n}$, $A\in\mathbb{R}^{m\times n}$, $a\in\mathbb{R}^{k}$, $b\in\mathbb{R}^{m}$ and $m,n,k,r\in\mathbb{N}$. (8) is called the cost or loss function, (9) represents the quadratic constraints, (10) are the linear constraints, and (11) enforces the integrality constraints for the last $r$ components of the vector of decision variables $x$.
Mixed-integer quadratic programs have been shown to be NP-hard [20], which often leads to an exhaustive demand for computational resources. The algorithms used to solve MIQP are typically branch-and-bound or cutting plane [16, 6, 54, 32]. Both of these algorithmic treatments are often employed together, often with the addition of a presolving step, the use of heuristics, and parallelism. The aforementioned allows many modern solvers to solve even large problems despite the NP-hardness. Some of these solvers are open source (like SCIP and GLPK), and others are commercial (GUROBI and CPLEX). The powerful infrastructure present in these solvers can be made use of together with additional problem-specific modifications to deliver high-quality solutions.
Due to the exhaustive nature of the algorithms mentioned in the previous paragraph, global convergence is guaranteed [3]. Furthermore, convergence to the global solution may be tracked and the error estimated by computing the dual problem of (811). The dual of the problem is then used to compute the so-called MIP GAP as follows
MIP $\text{MIP GAP}=\frac{\left|J\left(x^{*}\right)-J_{\text{dual}}\left(y^{*}\right)\right|}{\left|J\left(x^{*}\right)\right|},$ (12)
where $x^{*}$ and $y^{*}$ are the current best solutions of the primal and dual problems, respectively, and $J$ and $J^{*}$ are the cost functions of the primal and dual problems, respectively. The MIP GAP ensures that we can assess the quality of the minimization during solution time and terminate the computation when the result is good enough (small enough MIP GAP). Furthermore, if the gap reaches 0 at any point, we are sure that the current solution is a global optimum.